(dy)/(dx) = - x/(2y)


1
Expert's answer
2022-05-13T05:38:17-0400

Question

"\frac {dy} {dx}=-\frac {x} {2y}"

Solution.

We have:

"\frac {dy} {dx}=-\frac {x} {2y}"

"2y\,dy=-x\,dx"

"\int {2y} \ dy=\int {-x}\ dx"

"2\frac {y^2} {2}=-\frac {x^2} {2} + C"

"y^2=-\frac {x^2} {2} + C"

"y=\pm \sqrt{-\frac {x^2} {2} + C}"

"y^2=-\frac {x^2} {2} + C" is the general integral in an implicit form.